# SafeGen: Mitigating Unsafe Content Generation in Text-to-Image Models

### v1

@@ 6.1.1 NSFW Content Reduction
Table II: [RQ1-CLIP] Performance of SafeGen on reducing text-to-image alignment against different adversarial prompts compared with eight baseline methods.
@@
External Censorship & Filter SD-V2.1 20.30 19.19 21.75 23.90 24.60 23.66 24.02 24.08 24.81 22.21
@@
Internal Text-dependent ESD 19.89 18.12 21.16 24.59 24.04 24.11 24.59 24.72 25.94 23.79
@@
Different NSFW body parts.    Figure 10 shows the results of SafeGen and baseline methods in reducing the generation of various exposed body parts, e.g., M-Breasts or F-Breasts, on the NSFW-56k dataset, where ‘M’ stands for male and ‘F’ stands for female. SafeGen achieves a 99.1% NRR for total exposed body parts, while the others are less effective on some body parts. For example, SLD (Strong) exhibits a 22.2% NRR for buttocks, and SD-V2.1 has a 26.0% NRR for belly, which indicates their limitation on undefined or unseen NSFW concepts. Moreover, a -16.9% NRR on buttocks caused by SLD (Weak) suggests some safe measures can unintentionally steer the denoising diffusion process towards unsafe regions. Due to the page limitation, we display the removal results on other three datasets in Appendix E (Figure 12).
6.1.2 NSFW Text-to-image Alignment Reduction
Table II shows the results of SafeGen and baselines in reducing the text-to-image alignment among different adversarial prompts, rendering findings from two perspectives:
Overall effectiveness.    SafeGen outperforms all baselines in reducing the text-to-image alignment across all adversarial prompt datasets. We make two observations. Firstly, SafeGen consistently achieves the lowest CLIP scores compared with baselines, successfully severing the association between sexually explicit text information and visual representations. Notably, SafeGen demonstrates a minimal CLIP score variation of 2.67, whereas the others exhibit more significant fluctuations, e.g., ESD ranging from 18.12 to 24.59 (6.47) and SLD (Weak) ranging from 20.50 to 26.45 (5.95). This suggests the ability of SafeGen to maintain stable performance against varying adversarial prompts. Secondly, the NSFW-56k dataset serves as a good benchmark for assessing the effectiveness of anti-NSFW mitigation. Across the SneakyPrompt-N, SneakyPrompt-P, and I2P-Sexual datasets, the average CLIP score among all methods is 19.75 with a standard deviation of 1.74. In contrast, for the NSFW-56k dataset, the average CLIP score is higher at 23.50, with a larger standard deviation of 3.03. This comparison highlights the increased difficulty of the NSFW-56k dataset, characterized by a higher average score (indicating more sexually explicit generation by the models) and a greater standard deviation (indicating more instability of the method). Hence, the NSFW-56k provides a more distinct basis for evaluating the effectiveness of countermeasures.
@@ 6.1.2 NSFW Text-to-image Alignment Reduction
Table III: [RQ2] Performance of SafeGen in preserving the benign generation on COCO-25k prompts and comparison with baselines.
@@
CLIP Score normal-$\uparrow$ LPIPS Score normal-$\downarrow$ FID-25k normal-$\downarrow$
@@
Table III: [RQ2] Performance of SafeGen in preserving the benign generation on COCO-25k prompts and comparison with baselines.
6.2 RQ2: Benign Generation Preservation
We compare SafeGen with seven baselines in the ability to preserve the benign generation, as shown in Table III. We exclude the safety filter in this research question since it does not affect benign image generation as an external plug-in. We use COCO-25k as a reference dataset, which contains 5,000 benign images with 25,000 prompts, i.e., 5 annotated prompts for each image. We generate one image for each prompt. For each generated image, the CLIP score is calculated with its corresponding prompt, and we report the average score on all generated images. The LPIPS score is calculated individually between the generated and referenced images. The FID score is calculated between the set of generated images and the set of referenced images.
We present three key observations. Firstly, SafeGen achieves a CLIP score on par with the original SD, indicating its ability to preserve benign text-to-image preservation without degradation. In contrast, text-dependent methods with reasonable anti-NSFW levels such as ESD, and SLD (Max/Strong) have lower benign CLIP scores (ranging from 23.03 to 23.97), averaging 0.83 lower than SafeGen, which suggests a potential compromise in content alignment. Secondly, SafeGen’s LPIPS score and FID score are aligned with the original SD without decrease, which means SafeGen is capable of generating high-fidelity benign imagery. As a comparison, while text-dependent methods yield similar LPIPS scores, they show a higher average FID-25k gap of 3.0 than 20.31 of SafeGen, suggesting a potential negative impact on accurately reflecting textual descriptions in benign generation. Thirdly, the comparison of generated images between SafeGen and existing methods shown in Figure 13 suggests that SafeGen’s superior performance in human evaluation. It well maintains the images’ original style and overall layout of the original SD. While ESD obtains comparable performance in objective metrics like LPIPS and FID, it obviously affects the overall content and quality.
The reason is that the text-dependent methods erase or modify some NSFW concepts, e.g., nudity, in the SD model. Such modifications often pertain to human-related content, which is integral to the image’s context. As a result, altering these aspects can lead to a misalignment between the text and image, and also affect the model’s overall fidelity, especially for human-related objects.
6.3 RQ3: Performance Combined with Baselines
Table IV: [RQ3] Performance of SafeGen when combined with text-dependent mitigation methods in reducing NSFW generation while preserving benign generation.
@@ 6.3 RQ3: Performance Combined with Baselines
Ours (Vision-Only) 91.7% 17.79 0.805 24.33
Ours+SLD (Weak) 94.7% 17.84 0.787 24.33
Ours+SLD (Medium) 95.6% 17.16 0.790 23.77
Ours+SLD (Strong) 97.1% 16.83 0.794 23.29
Ours+SLD (Max) 97.8% 16.75 0.802 22.85
Ours+ESD 95.7% 19.93 0.795 24.12
Original SD - 21.77 0.782 24.56
@@
ESD 80.3% 19.89 0.788 23.77
SLD (Max) 72.7% 18.63 0.801 23.03
Safety Filter 67.9% 19.01 / /
Table IV: [RQ3] Performance of SafeGen when combined with text-dependent mitigation methods in reducing NSFW generation while preserving benign generation.
Table IV shows the results of the performance of SafeGen when combined with baselines. We evaluate the combination with text-dependent baselines, i.e., ESD and different variants of SLD on both NSFW mitigation and benign preservation. We skip the safety filter baseline since it has no impact on benign generation. We employ SneakyPrompt-N for testing NSFW mitigation and COCO-25k for testing benign preservation. In each dataset, we randomly select 200 prompts, and then generate three images per prompt using different random seeds. Our findings reveal that SafeGen, with only self-attention layers adjustment, alone outperforms all baselines in terms of NSFW removal with average 21.7% NRR improvement, while retaining high-fidelity benign generation with comparable CLIP score. In addition, the integration with other text-dependent techniques demonstrate SafeGen significantly aids baselines in reducing NSFW content generation, realizing a remarkable 27.8% NRR enhancement.
From the perspective of NSFW content mitigation, SafeGen + SLD (Max) achieves the highest NRR at 97.8% and the lowest CLIP score at 16.75, indicating its effectiveness in mitigating exposed body parts generation and deviating the resulting images from adversarial prompts. On the other hand, from the perspective of benign generation mitigation, SafeGen + SLD (Weak) has the lowest LPIPS score at 0.787 and the highest CLIP score at 24.33, which suggests it preserves the visual fidelity of benign images well.
This observation suggests a trade-off between NSFW generation mitigation and benign generation preservation. While SafeGen + SLD (Max) is most effective in NSFW removal, it slightly compromises image fidelity as indicated by a higher LPIPS score. Conversely, SafeGen + SLD (Weak) preserves benign image fidelity better but does not perform as well in NSFW removal as the former. Thus, the choice of method depends on the specific requirements of the task, i.e., whether the priority is to maximize NSFW content removal or to preserve the fidelity of benign images.
Table V: [RQ4-1] Performance of SafeGen across different hyperparameters $\lambda_{m}$ and $\lambda_{p}$.
@@
$\lambda_{m}:0.1,\lambda_{p}:0.9$ 95.2% 17.85 0.789 24.60
$\lambda_{m}:0.2,\lambda_{p}:0.8$ 96.1% 17.64 0.792 24.21
$\lambda_{m}:0.3,\lambda_{p}:0.7$ 98.6% 17.12 0.814 24.17
$\lambda_{m}:0.4,\lambda_{p}:0.6$ 92.3% 17.91 0.822 24.12
$\lambda_{m}:0.5,\lambda_{p}:0.5$ 97.6% 17.30 0.839 23.87
Table V: [RQ4-1] Performance of SafeGen across different hyperparameters $\lambda_{m}$ and $\lambda_{p}$.
6.4 RQ4: Exploration on Hyperparameters
This subsection explores the impact of different hyperparameters on SafeGen. Specifically, we examine $\lambda_{m}$ and $\lambda_{p}$, which are responsible for mitigating model non-compliance while preserving the model’s ability to generate benign images. Moreover, we investigate the influence of distinct diffusion schedulers and varied diffusion steps in the inference stage. We set the number of generated images per prompt to one, considering the computational overhead in image generation by T2I models during extensive parameter comparison. We ensure the reliability of our findings through statistical analysis across two datasets.
6.4.1 Different Hyperparameters of Loss Weights
Table V presents the performance of SafeGen on both NSFW removal and benign preservation by varying the loss weights $\lambda_{m}$ and $\lambda_{p}$. We observe an overall upward trend of NRRs and downward trend of CLIP scores by gradually increasing $\lambda_{m}$ and decreasing $\lambda_{p}$, denoting a larger $\lambda_{m}$ and a smaller $\lambda_{p}$ benefits NSFW content removal and suppresses the adversarial text-to-image alignment. The average NRR and CLIP score under different $\lambda_{m,p}$ combinations are up to 95.9% and down to 17.57, with negligible variance, suggesting SafeGen can well mitigate NSFW concepts with a wide parameter space. Prompted by benign texts from the COCO-25k dataset, results demonstrate a lower $\lambda_{m}$ and higher $\lambda_{p}$ can ensure that SafeGen yields high-fidelity images. The optimal LPIPS score is down to 0.789 and the best CLIP score reaches 24.60, even slightly surpasses the original SD’s performance with 24.56. This experiment guides us in selecting the optimal hyperparameter combination, i.e., $\lambda_{m}$: 0.1 and $\lambda_{p}$: 0.9, that excels in defense efficacy while preserving the fidelity of benign image generation from a systematic performance perspective.

### v2

@@ 6.1.2. Explicit Text-to-image Alignment Reduction
Table 2. [RQ1-CLIP] Performance of SafeGen on reducing text-to-image alignment against different adversarial prompts compared with eight baseline methods.
@@
Censorship & Filter (External) SD-V2.1 20.30 19.19 21.75 23.90 24.60 23.66 24.02 24.08 24.81 22.21
@@
Text-dependent (Internal) ESD 19.89 18.12 21.16 24.59 24.04 24.11 24.59 24.72 25.94 23.79
@@
Overall effectiveness. SafeGen outperforms all baselines in reducing the text-to-image alignment across all adversarial prompt datasets. We make two observations. Firstly, SafeGen consistently achieves the lowest CLIP scores compared with baselines, successfully severing the association between sexually explicit text information and visual representations. Notably, SafeGen demonstrates a minimal CLIP score variation of 2.67, whereas the others exhibit more significant fluctuations, e.g., ESD ranging from 18.12 to 24.59 (6.47) and SLD (Weak) ranging from 20.50 to 26.45 (5.95). This suggests the ability of SafeGen to maintain stable performance against varying adversarial prompts. Secondly, the NSFW-56k dataset serves as a good benchmark for assessing the effectiveness of sexually explicit mitigation. Across the SneakyPrompt-N, SneakyPrompt-P, and I2P-Sexual datasets, the average CLIP score among all methods is 19.75 with a standard deviation of 1.74. In contrast, for the NSFW-56k dataset, the average CLIP score is higher at 23.50, with a larger standard deviation of 3.03. This comparison highlights the increased difficulty of the NSFW-56k dataset, characterized by a higher average score (indicating more sexually explicit generation by the models) and a greater standard deviation (indicating more instability of the method). Hence, the NSFW-56k provides a more distinct basis for evaluating the effectiveness of countermeasures.
@@
Table 3. [RQ2] Performance of SafeGen in preserving the benign generation on COCO-25k prompts and comparison with baselines.
@@
CLIP Score $\uparrow$ LPIPS Score $\downarrow$ FID-25k $\downarrow$
@@
6.2. RQ2: Benign Generation Preservation
We compare SafeGen with seven baselines in the ability to preserve the benign generation, as shown in Table 3. We exclude the safety filter in this research question since it does not affect benign image generation as an external plug-in. We use COCO-25k as a reference dataset, which contains 5,000 benign images with 25,000 prompts, i.e., 5 annotated prompts for each image. We generate one image for each prompt. For each generated image, the CLIP score is calculated with its corresponding prompt, and we report the average score on all generated images. The LPIPS score is calculated individually between the generated and referenced images. The FID score is calculated between the set of generated images and the set of referenced images.
We present three key observations. Firstly, SafeGen achieves a CLIP score on par with the original SD, indicating its ability to preserve benign text-to-image preservation without degradation. In contrast, text-dependent methods with reasonable anti-sexually-explicit levels such as ESD, and SLD (Max/Strong) have lower benign CLIP scores (ranging from 23.03 to 23.97), averaging 0.83 lower than SafeGen, which suggests a potential compromise in content alignment. Secondly, SafeGen’s LPIPS score and FID score are aligned with the original SD without decrease, which means SafeGen is capable of generating high-fidelity benign imagery. As a comparison, while text-dependent methods yield similar LPIPS scores, they show a higher average FID-25k gap of 3.0 than 20.31 of SafeGen, suggesting a potential negative impact on accurately reflecting textual descriptions in benign generation. Thirdly, the comparison of generated images between SafeGen and existing methods shown in Figure 17 suggests that SafeGen’s superior performance in human evaluation. It well maintains the images’ original style and overall layout of the original SD. While ESD obtains comparable performance in objective metrics like LPIPS and FID, it obviously affects the overall content and quality.
The reason is that the text-dependent methods erase or modify some NSFW concepts (e.g., nudity, sexual), in the SD model. Such modifications often pertain to human-related content, which is integral to the image’s context. As a result, altering these aspects can lead to a misalignment between the text and image, and also affect the model’s overall fidelity, especially for human-related objects.
6.3. RQ3: Performance Combined with Baselines
@@
Table 4. [RQ3] Performance of SafeGen when combined with text-dependent mitigation methods in reducing sexually explicit generation while preserving benign generation.
@@
Ours (Vision-Only) 92.8% 17.79 0.805 24.33
Ours+SLD (Weak) 95.5% 17.84 0.787 24.33
Ours+SLD (Medium) 96.0% 17.16 0.790 23.77
Ours+SLD (Strong) 97.3% 16.83 0.794 23.29
Ours+SLD (Max) 98.2% 16.75 0.802 22.85
Ours+ESD 96.0% 19.93 0.795 24.12
Original SD 0% 21.77 0.782 24.56
@@
ESD 84.2% 19.89 0.788 23.77
SLD (Max) 81.8% 18.63 0.801 23.03
Safety Filter 71.2% 19.01 / /
Table 4 shows the results of the performance of SafeGen when combined with baselines. We evaluate the combination with text-dependent baselines, i.e., ESD and different variants of SLD on both nudity mitigation and benign preservation. We skip the safety filter baseline since it has no impact on benign generation. We employ SneakyPrompt-N for testing nudity mitigation and COCO-25k for testing benign preservation. In each dataset, we randomly select 200 prompts, and then generate three images per prompt using different random seeds. Our findings reveal that SafeGen, with only self-attention layers adjustment, alone outperforms all baselines in terms of nudity removal with average 21.7% NRR improvement, while retaining high-fidelity benign generation with comparable CLIP score. In addition, the integration with other text-dependent techniques demonstrate SafeGen significantly aids baselines in reducing sexually explicit content generation, realizing a remarkable 27.8% NRR enhancement.
From the perspective of nudity content mitigation, SafeGen + SLD (Max) achieves the highest NRR at 97.8% and the lowest CLIP score at 16.75, indicating its effectiveness in mitigating exposed body parts generation and deviating the resulting images from adversarial prompts. On the other hand, from the perspective of benign generation mitigation, SafeGen + SLD (Weak) has the lowest LPIPS score at 0.787 and the highest CLIP score at 24.33, which suggests it preserves the visual fidelity of benign images well.
This observation suggests a trade-off between unsafe generation mitigation and benign generation preservation. While SafeGen + SLD (Max) is most effective in nudity removal, it slightly compromises image fidelity as indicated by a higher LPIPS score. Conversely, SafeGen + SLD (Weak) preserves benign image fidelity better but does not perform as well in nudity removal as the former. Thus, the choice of method depends on the specific requirements of the task, i.e., whether the priority is to maximize sexually explicit content removal or to preserve the fidelity of benign images.
@@
Table 5. [RQ4-1] Performance of SafeGen across different hyperparameters $\lambda_{m}$ and $\lambda_{p}$.
@@
$\lambda_{m}:0.1,\lambda_{p}:0.9$ 99.0% 17.85 0.789 24.60
$\lambda_{m}:0.2,\lambda_{p}:0.8$ 97.6% 17.64 0.792 24.21
$\lambda_{m}:0.3,\lambda_{p}:0.7$ 99.0% 17.12 0.814 24.17
$\lambda_{m}:0.4,\lambda_{p}:0.6$ 93.7% 17.91 0.822 24.12
$\lambda_{m}:0.5,\lambda_{p}:0.5$ 98.6% 17.30 0.839 23.87
6.4. RQ4: Exploration on Hyperparameters
This subsection explores the impact of different hyperparameters on SafeGen. Specifically, we examine $\lambda_{m}$ and $\lambda_{p}$, which are responsible for mitigating model non-compliance while preserving the model’s ability to generate benign images. Moreover, we investigate the impact of training data selection in the model editing stage, as well as distinct diffusion schedulers and varied diffusion steps in the inference stage. We set the number of generated images per prompt to one, considering the computational overhead in image generation by T2I models during extensive parameter comparison. We ensure the reliability of our findings through statistical analysis across two datasets.
6.4.1. Different Hyperparameters of Loss Weights
Table 5 presents the performance of SafeGen on both nudity removal and benign preservation by varying the loss weights $\lambda_{m}$ and $\lambda_{p}$. We observe an overall upward trend of NRRs and downward trend of CLIP scores by gradually increasing $\lambda_{m}$ and decreasing $\lambda_{p}$, denoting a larger $\lambda_{m}$ and a smaller $\lambda_{p}$ benefits nudity content removal and suppresses the adversarial text-to-image alignment. The average NRR and CLIP score under different $\lambda_{m,p}$ combinations are up to 95.9% and down to 17.57, with negligible variance, suggesting SafeGen can well mitigate nudity concepts with a wide parameter space. Prompted by benign texts from the COCO-25k dataset, results demonstrate a lower $\lambda_{m}$ and higher $\lambda_{p}$ can ensure that SafeGen yields high-fidelity images. The optimal LPIPS score is down to 0.789 and the best CLIP score reaches 24.60, even slightly surpasses the original SD’s performance with 24.56. This experiment guides us in selecting the optimal hyperparameter combination, i.e., $\lambda_{m}$: 0.1 and $\lambda_{p}$: 0.9, that excels in defense efficacy while preserving the fidelity of benign image generation from a systematic performance perspective.